OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work. The pictures are in the comments. 1. 2x^4(4x^2 + 3x + 1) 2. (4x – 3)(2x^2 – 7x + 1) 3. (x^2 + 4x – 3)(2x^2 + x + 6) 4. Write a simplified polynomial expression to represent the area of the rectangle below. 5. Write a simplified polynomial expression to represent the area of the square tile, shown below.

OpenStudy (anonymous):

The pictures are for 4 and 5

OpenStudy (anonymous):

the top is for number 4 the bottom is for number 5

OpenStudy (anonymous):

please help!

OpenStudy (mysesshou):

Since I can't get my work done, I'll try a little bit of yours. :) This is a great site. http://www.wolframalpha.com/input/?i=simplify++2x%5E4%284x%5E2+%2B+3x+%2B+1%29+ ok. now I'll look at it more.

OpenStudy (anonymous):

i have tried that site before it isn't always accurate and thank you so much! i owe you one!

OpenStudy (mysesshou):

1. 2x^4(4x^2 + 3x + 1) multiply the \[2x ^{4}\] through to the others. \[(2x ^{4}*4x ^{2})+(2x^{4}*3x)+(2x^{4}*1) \] \[=8x^{6}+6x^{5}+2x^{4}\] Ok. I multiplied #1 myself. The site was right for that one. :)

OpenStudy (mysesshou):

exponents next to each other, like x^2*x^3 they add. to give x^5. exponents multiply if they are to the power of each other, like (x^2)^3, gives x^6.

OpenStudy (mysesshou):

(4x – 3)(2x^2 – 7x + 1) \[4x*2x^{2} - 4x*7x +4x*1 -3*2x^{2} +3*7x -3*1\] Note the +3*7x, since the negative*negative=positive. Multiply out the terms: \[8x^{3}-28x^{2}+4x-6x^{2}+21x-3\] wow, this is more difficult on the computer than by hand. :) Combine like terms (simplify): \[8x^{3}-34x^{2}+25x-3\] My computer is so slow :( Double check that my sol is "standard form"...

OpenStudy (anonymous):

it is :)

OpenStudy (mysesshou):

The same is done for #3. do you still need help w that one too?

OpenStudy (anonymous):

yes my brain is fried if it isnt any trouble

OpenStudy (mysesshou):

OK. Mine too. that is why i'm avoiding my Q's ok. give a sec

OpenStudy (mysesshou):

(x^2 + 4x – 3)(2x^2 + x + 6) Let's take the first () multiplied through the second () : x^2*(2x^2 + x + 6) + 4x*(2x^2 + x + 6) - 3*(2x^2 + x + 6) \[(x^2*2x^2 + x^2*x + x^2*6) + (4x*2x^2 + 4x*x + 4x*6) - (3*2x^2 + 3*x + 3*6) \] \[(2x^4 + x^3 + 6x^2) + (8x^3 + 4x^2 + 24x) - (6x^2 + 3x + 18) \] remove the (), but take into account that minus, so it is like multiplying (-1) through the whole last () set. \[2x^4 + x^3 + 6x^2 +8x^3 + 4x^2 + 24x - 6x^2 - 3x - 18 \] combine like terms: \[2x^4 + 9x^3 + 4x^2 + 21x - 18 \] and it matches that site, yay or me. :) typing really is more difficult. :)

OpenStudy (mysesshou):

*yay for me... typo

OpenStudy (anonymous):

thank you i can do the rest!

OpenStudy (mysesshou):

ok! good luck!

OpenStudy (anonymous):

if you ever need anyhting just message me!

OpenStudy (anonymous):

can you help with one more question?

OpenStudy (mysesshou):

I need college probability and linear algebra :) I posted a Q but no one has gotten to it yet. No worries though.

OpenStudy (anonymous):

oh let me see if i can help, not good at the stuff though lol

OpenStudy (mysesshou):

You have two matrices of the same size. Compare rank(A+B) and rank(A)+rank(B) and prove.

OpenStudy (anonymous):

ahhh nope can't do that lol

OpenStudy (anonymous):

thought you meant something different

OpenStudy (mysesshou):

no worries. i didnt take the class, that is why i don't know how to do it. :) ok, let me read your other link

OpenStudy (mysesshou):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x - 6) and (x - 6)(x - 6) Ok, the hoblos reply is right. \[(x+6)(x-6) = x^2 -6x +6x -36\] which simplifies by the -6x and +6x cancelling. \[x^2-36\] Yes, this is difference of two squares, http://en.wikipedia.org/wiki/Difference_of_two_squares I suppose you could compare that they both give the x^2 and the 36, but due to the sign differences the 36 is positive for the one below and negative for the one above. Also, due to the sign difference in the one above, the -6x cancels the +6x. I'm not sure what else to add. I'll have to think a bit. \[(x - 6)(x - 6) = (x-6)^2 = x^2 -6x-6x +36=x^2 -12x +36\]

OpenStudy (anonymous):

OKAY THANK YOU!

OpenStudy (anonymous):

oops caps locks

OpenStudy (mysesshou):

welcome!